# TRACE: Learning to Compute on Graphs

### v1

@@ 3.3 Function Shift Learning on Predictive Task
By ignoring the dependencies among input, i.e. by assuming they are independent, we can derive the local function, which is formally stated in Definition 2. Although this approximation is computationally efficient with an $O(1)$ complexity, it fails to capture the true function of the circuit.
@@
Building on these properties, we propose to learn the function shift: $y_{\phi}^{FSL}=y_{\phi}^{global}-y_{\phi}^{local},$ which measures the discrepancy between the local function (Definition 2) and the true global function (Definition 1). This formulation decouples the global function into two components: a simple local function and a function shift. Rather than predicting the complex global function directly, our model is trained to predict only the function shift. This isolates the complex contextual effects caused by reconvergence, allowing the global function to be reconstructed by simply combining the predicted shift with the local function.
@@
$\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\Big[\big|\psi(x_{i})-y_{i}^{FSL}\big|\Big],$ (5)
where $x_{i}$ is the final embedding for node $i$ and $\psi(\cdot)$ is a 3-layer MLP regression head.
Inference Stage.
@@
7:   $\begin{aligned} \hat{y}_{v}^{global}\leftarrow\hat{y}_{v}^{FSL}+&\\ \underbrace{\phi_{v}(\hat{y}_{u_{1}}^{global},\dots,\hat{y}_{u_{|\mathcal{N}(v)|}}^{global})}_{\text{Local Function}}\end{aligned}$
8:  end for
9:end for
10:Return $\hat{\mathbf{y}}^{global}$ for all nodes
3.4 Contrastive Task
Contrastive learning is a standard self-supervised strategy for learning representations of circuit functionality (Wang et al., 2024; Fang et al., 2025b; c; Wu et al., 2025). The fundamental principle is to pull embeddings of functionally equivalent circuits closer together in the embedding space while pushing apart those of functionally different circuits. This process trains the model to identify and encode the discriminative features that define a circuit’s intrinsic properties, all without needing explicit labels. Following prior work, we form training instances for each circuit $\mathcal{G}$. A positive sample, $\mathcal{G}^{+}$, is created by applying a functionally equivalent transformation to $\mathcal{G}$. All other circuits within the same batch are treated as negative samples, $\mathcal{G}^{-}$. We then optimize the encoder using the InfoNCE loss (Oord et al., 2018):
$\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\;\mathcal{L}_{\text{InfoNCE}}(\mathcal{G},\mathcal{G}^{+},\mathcal{G}^{-}).$ (6)
The effect of this objective is to structure the embedding space such that functional equivalence corresponds to proximity, mapping similar circuits to nearby points while separating them from non-equivalent ones.
4 Experiment
4.1 Implement Details
Dataset
In this paper, we conduct experiments on three modalities: RTL, AIG and PM netlist. For RTL, we follow previous works (Fang et al., 2025b; c) and collect data from ITC (Corno et al., 2002) and OpenCores (Albrecht, 2005). For combinational AIGs, we use ForgeEDA (Shi et al., 2025a). For Sequential AIGs, we follow DeepSeq2 (Khan et al., 2025) and extract sub-circuits from ITC (Corno et al., 2002), OpenCores (Albrecht, 2005) and ISCAS’89 (Brglez et al., 1989). For PM netlist, we follow previous work (Shi et al., 2025b) and extract sub-circuits from ForgeEDA (Shi et al., 2025a) and DeepCircuitX (Li et al., 2025). More details are provided in Appendix C.
Evaluation Metrics
In this work, we evaluate our model on two types of tasks: contrastive and predictive. For the first type, the contrastive task, the goal is retrieval. Given a query circuit, the model must identify its functionally equivalent (positive) counterpart from a pool of $N$ candidate circuits. For our experiments, we set the pool size to $N=256$ for RTL and $N=1024$ for AIG and PM Netlists. We measure performance using the Recall@k (Rec@k) metric, reporting scores for $k\in\{1,5,10\}$. For the second type, the predictive tasks, we assess the model’s ability to determine node-level properties. We follow previous works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2025b) and perform logic-1 probability prediction, similarity prediction and transition probability prediction (See Appendix E). After encoding a circuit to produce node embeddings, these are used to predict a target value for each node. We evaluate the accuracy of these predictions using Mean Absolute Error (MAE) and the $R^{2}$ score.
Table 1: Comparison of contrastive task across various modalities(%).
@@ 4.1 Implement Details
4.2 Contrastive Tasks
RTL
On the RTL modality, traditional message-passing models such as GCN (Kipf, 2016), GIN (Xu et al., 2018), GAT (Veličković et al., 2017), and GraphSAGE (Hamilton et al., 2017) achieve moderate performance, with Rec@1 ranging from 82.9% to 86.5%. Graph Transformer variants like GraphGPS (Rampášek et al., 2022), SGFormer (Wu et al., 2023c), DIFFormer (Wu et al., 2023b), and the RTL-specialized CircuitEncoder (Fang et al., 2025c) show competitive scores, with Rec@1 ranging from 79.45% to 88.28%. In contrast, TRACE achieves a substantial improvement, pushing Rec@1 to 94.45% and Rec@10 to 99.89%, outperforming the second-best method by 6.17% and 2.94% respectively.
AIGs
For AIGs, the performance gap between baselines and TRACE becomes even more pronounced. While traditional message-passing networks perform decently (with Rec@1 from 83.01% to 88.55%) and AIG-specialized architectures like FGNN2 (Wang et al., 2024) reach 88.7%, Graph Transformer models generally struggle. For instance, performance of SGFormer and DIFFormer drop to as low as 15.4% and 37.03% Rec@1 respectively, highlighting the drawback of methods that destroy the explicit hierarchy in a computational graph. GraphGPS even suffers from an Out-of-Memory (OOM) error due to its dense attention mechanism. Our method, however, consistently outperforms all baselines, reaching 92.68% at Rec@1 and 99.51% at Rec@10.
PM Netlists
The Netlist modality presents the most challenging benchmark, where message-passing models show varying performance. For instance, GraphSAGE and the netlist-specialized DeepCell (Shi et al., 2025b) achieve strong results (Rec@1 around 80–86%), while others like GCN and GAT show a significant performance drop. Graph Transformer models also suffer from severe performance degradation, with Rec@1 scores below 50% for GraphGPS, SGFormer, and DIFFormer. Remarkably, TRACE delivers consistent and superior performance, attaining 90.81% Rec@1 and 99.44% at Rec@10, outperforming both families of baselines by a large margin.
In summary, TRACE consistently outperforms both message-passing and Transformer-based baselines across all three modalities. Furthermore, its stable performance across these diverse graph types highlights its strong capacity for generalization.
Table 2: Comparison of predictive tasks on combinational and sequential AIGs.
@@ 4.2 Contrastive Tasks
4.3 Predictive Tasks
Combinational AIGs
On combinational AIGs, message-passing baselines such as GCN, GIN, GAT, GraphSAGE, PolarGate (Liu et al., 2024) and MGVGA (Wu et al., 2025) yield poor performance, with $R^{2}$ values ranging from 0.493 to 0.675 for logic-1 probability prediction and from 0.021 to 0.445 for similarity prediction. Graph Transformer baselines either suffer from OOM or achieve limited performance. AIG-specialized architectures, like DeepGate2 (Shi et al., 2023), DeepGate4 (Zheng et al., 2025) improves results with $R^{2}$ around 0.98 for logic-1 probability and around 0.50 for similarity. In contrast, TRACE delivers the best overall performance, achieving an $R^{2}$ of 0.989 with MAE 0.015 for logic-1 probability and significantly outperforming all baselines in similarity prediction with $R^{2}$ of 0.633 and MAE of 0.055.
Sequential AIGs
Sequential circuits pose additional challenges due to temporal dependencies. Our method again demonstrates substantial gains, reaching an $R^{2}$ of 0.997 with MAE 0.009 for logic-1 probability, and 0.976 with MAE 0.005 for transition probability, outperforming the second-best method, DeepSeq2 (Khan et al., 2025), by 0.018 in R2 on logic-1 probability prediction and 0.068 in R2 on transition probability prediction.
PM Netlists
The predictive task on PM netlists further validates the generalization ability of TRACE. Message-passing methods achieve moderate performance, with GraphSAGE and DeepCell reaching $R^{2}$ above 0.94. Graph Transformer models show mixed results, with SGFormer performing relatively well ($R^{2}=0.918$) but DIFFormer dropping to 0.696. TRACE clearly surpasses all baselines, achieving an $R^{2}$ of 0.994 and a minimal MAE of 0.013. This consistent superiority across modalities emphasizes the adaptability of our approach to different circuit representations and predictive objectives.
Overall, across all predictive tasks, TRACE not only surpasses both message-passing and Transformer-based models, but also approaches near-perfect accuracy, demonstrating its capacity to generalize across combinational, sequential, and physical design graph domains.
Table 3: Comparison of predictive task on PM netlists.
Model R2 MAE
@@ 4.3 Predictive Tasks
Message Passing Neural Network
GCN 0.718 0.112
GraphSAGE 0.946 0.048
GAT 0.902 0.059
GIN 0.734 0.102
DeepCell 0.942 0.053
Graph Transformer
GraphGPS 0.846 0.083
SGFormer 0.918 0.056
DIFFormer 0.696 0.141
TRACE 0.994 0.013
4.4 Ablation Study
Table 4: Ablation study on function shift learning (FSL).
Setting PM Netlist AIG
Logic-1 Probability Logic-1 Probability Similarity Prediction
R2 MAE R2 MAE R2 MAE
TRACE w/o FSL 0.985 0.036 0.980 0.024 0.500 0.066
TRACE 0.994 0.013 0.989 0.015 0.533 0.055
We conducted an ablation study to quantify the contribution of our proposed Function Shift Learning (FSL) component, which is designed to help our model adapt to different graph types and their unique functional distributions. The results, summarized in Table 4, demonstrate that FSL is a crucial element for achieving high performance. The model with FSL consistently outperforms its ablated counterpart (TRACE w/o FSL) across all tasks and datasets. For Logic-1 Probability prediction on PM Netlists, the addition of FSL significantly reduces the MAE from 0.036 to 0.013 and increases the R2 score from 0.985 to 0.994, indicating more accurate predictions. As for AIGs, FSL improves the MAE by 0.009 on logic-1 probability prediction and 0.011 on similarity prediction. This confirms that the FSL component is essential for our model’s ability to learn functional representations, leading to superior performance on diverse computational graphs.
5 Conclusion
@@
In this work, we introduced TRACE, a new paradigm for learning on computational graphs that addresses the architectural limitations of conventional MPNNs and Transformers. By employing a novel Hierarchical Transformer and a function shift learning objective, TRACE directly models the position-aware, hierarchical nature of computation. Our extensive experiments on electronic circuits demonstrate that TRACE substantially outperforms all prior architectures, establishing a new state of the art. This work provides a proof of principle for a more architecturally sound approach to learning on computational graphs, offering a powerful framework with potential applications across various domains.
Appendix A The Use of Large Language Models
In the preparation of this paper, we utilized a large language model (LLM) as an assistive tool to enhance the quality of our writing and presentation. The LLM’s role was strictly confined to refining the manuscript’s writing and formatting, without generating any core scientific content or data.
Appendix B Computation Graph
Register Transfer Level(RTL)
RTL design is a hardware abstraction used in the early stages of digital chip design, serving as the bridge between high-level behavioral descriptions and gate-level implementations. To be specific, RTL code captures how data moves between registers (i.e., sequential registers) and how logic gates operate on that data within each clock cycle (i.e., combinational logic). Essentially, an RTL design can be viewed as a directed graph. We first convert the RTL code in Hardware Description Language (HDL) format into an abstract syntax tree (AST) and then extract the graph structure based on this tree. In this graph, nodes represent word-level register signals and various operators (e.g., And, Add, Equal, Mux), while the wires in the HDL code form the edges that denote the paths of data flow.
And-Inverter Graph(AIG)
In our work, we use combinational AIG and sequential AIG, which are widely used for circuit analysis, optimization, and formal verification because of their compact and canonical representation of Boolean functions.
Combinational AIG is a directed acyclic graph (DAG) composed of three basic elements: Primary Input(PI), AND gate and NOT gate. Since any Boolean logic expression can be constructed using only AND and NOT operations, AIG provides a universal and efficient representation. For example, a simple logic expression $\lnot A\land B$ can be built as a DAG with 2 PIs(A and B), one NOT gate and one AND gate. The edges are $[(A,NOT),(NOT,AND),(B,AND)]$. The AND gate with no outgoing edges represents the circuit’s final output in this DAG.
Sequential AIG extends this by introducing registers as an additional node type. These registers can capture the circuit’s state at each clock cycle, enabling sequential AIG to represent more complex circuits with memory functionality, such as finite state machines.
Post-Mapping Netlist
A Post-Mapping Netlist is a gate-level representation obtained after logic synthesis and technology mapping, where the circuit is expressed using standard cells from a target technology library and optimized for timing, area, and power. Unlike AIGs, which represent circuits as abstract DAGs of AND and NOT gates (and optionally registers for sequential circuits) focusing on low-level function, post-mapping netlists capture high-level implementation details, including specific gate types and connectivity imposed by the target library. Consequently, the node types and structures can differ significantly from those in AIGs, which is why we treat AIGs and Post-Mapping (PM) netlists as distinct modalities in this paper.
Appendix C Dataset Details
We summarize the statistics of the datasets used in both contrastive tasks (Table 5) and predictive tasks (Table 6). For each dataset, we report the number of graphs (#Graphs), the number of nodes (#Nodes), the number of edges (#Edges), and the maximum depth of a graph (Depth). For #Nodes, #Edges, and Depth, we provide the minimum, average, and maximum values, denoted as $\{\text{min},\text{avg},\text{max}\}$.
Table 5: Dataset statistics for contrastive tasks with $\{min.,avg.,max.\}$.
RTL AIG PM Netlist
#Graphs 1138 67706 67728
#Nodes { 9.0, 102.6, 1987.0} {22.0, 162.2, 2127.0} {16.0, 88.6, 1192.0}
#Edges {10.0, 278.3, 2645.0 } {22.0, 176.0, 2275.0} {20.0, 110.5, 1361.0}
Depth {2.0, 8.9, 27.0} {6.0, 16.7, 29.0} {2.0, 5.9, 12.0}
Table 6: Dataset statistics for predictive tasks with $\{min.,avg.,max.\}$.
Com. AIG Seq. AIG PM Netlist
#Graphs 9800 10007 83042
#Nodes {10.0, 2324.9, 45409.0} {23.0, 236.5, 1281.0} {12.0, 668.1, 4783.0}
#Edges {12.0, 3268.57, 68676.0} {21.0, 260.5, 1915.0} {11.0, 1113.9, 8914.0}
Depth {4.0, 49.1, 2657.0} {4.0, 21.0, 102.0} {1.0, 16.0, 260.0}
Appendix D Degree Distribution
Figure 4: In-degree distribution. The x-axis represents the node in-degree and the y-axis represents the frequency (number of nodes).
In this section, we compare the in-degree properties of computational graphs and general graphs, and analyze the corresponding padding overhead induced by our method. Specifically, we study three types of computational graphs (RTL, AIG, and PM netlist) and compare them with citation graphs, a representative class of DAGs. For the latter, we use the Cora, CiteSeer, and PubMed datasets from Yang et al. (2016).
As shown in Figure 4, the in-degree of a node in a computational graph is largely determined by its operator type, leading to a stable and bounded distribution across RTL, AIG, and PM netlist. In contrast, general DAGs such as citation networks exhibit long-tailed in-degree distributions. This discrepancy is critical for our method (Section 3.2), as sequence length is padded according to node in-degree. Consequently, long-tailed distributions introduce significant redundancy for general DAGs, while computational graphs remain more compact.
We quantify the redundancy using the following metric:
$\text{padding\_overhead}=\frac{n\times\max_{i}(d_{i})-\sum_{i}d_{i}}{n\times\max_{i}(d_{i})},$ (7)
where $n$ denotes the number of nodes in the graph and $d_{i}$ is the in-degree of node $i$.
Table 7: Padding Overhead Across Different Graphs
Metric Computational Graph General Graph
RTL AIG PM Netlist Cora CiteSeer PubMed
Padding Overhead 42.35% 16.29% 38.54% 97.10% 96.22% 96.80%
As summarized in Table 7, citation graphs suffer from severe padding overhead, ranging from 96.22% to 97.10%, which corresponds to nearly $20\times$ additional computation cost. In contrast, computational graphs exhibit much lower overhead, between 16.29% and 42.35%. These results underscore the structural advantage of computational graphs: their bounded in-degree leads to substantially reduced padding, thereby improving the efficiency of our proposed method.
Appendix E Predictive Tasks
E.1 Logic-1 Probability Prediction
Logic-1 probability prediction is a node-level regression task. In a digital circuit, the logic value of any node $i$ can be modeled as a binary random variable $x_{i}\sim\mathcal{B}(p_{i})$, where $p_{i}$ is the probability of the node being in the logic “1” state. This value, often referred to as the signal probability, is a crucial indicator of circuit function. The objective of this task is to predict the parameter $p_{i}$ for each node in the circuit, providing insight into its functional behavior.
E.2 Functional Similarity Prediction
The core objective of similarity prediction is to predict the functional similarity between a given pair of nodes. To establish the ground truth for this task, we first sample a fixed set of input patterns from the complete input space (exhaustive simulation would require testing all $2^{n}$ possible input combinations for $n$ inputs). We sample a set of node pairs $\mathcal{N}_{\text{pairs}}$. For each node $i$ in a selected pair, we generate a partial truth table $T_{i}$ by recording its state under this shared set of input patterns. The functional similarity $S_{(i,j)}$ for a pair of nodes $(i,j)\in\mathcal{N}_{\text{pairs}}$ is then calculated based on the normalized Hamming distance between their respective partial truth tables, $T_{i}$ and $T_{j}$:
$S_{(i,j)}=1-\frac{\text{HammingDistance}(T_{i},T_{j})}{\text{length}(T_{i})}$ (8)
This similarity score $S_{(i,j)}$ ranges from 0 to 1, where 1 indicates that the two nodes have identical outputs for all simulated input vectors, and 0 indicates they are completely dissimilar.
E.3 Transition Probability Prediction
To analyze the dynamic behavior of sequential circuits, random binary input sequences are applied to each primary input, and the circuit is simulated. For each input sequence, the output states $s_{i}(t)\in\{0,1\}$ of standard cells and the sequential outputs of registers are recorded. The number of transitions from 0 to 1 and from 1 to 0 for each cell or register $i$ is counted as $N_{0\to 1}^{(i)}$ and $N_{1\to 0}^{(i)}$, respectively. The transition probabilities are then defined as
$P_{0\to 1}^{(i)}=\frac{N_{0\to 1}^{(i)}}{N_{\text{total}}},\quad P_{1\to 0}^{(i)}=\frac{N_{1\to 0}^{(i)}}{N_{\text{total}}},$ (9)
where $N_{\text{total}}$ is the total number of input sequences.
Notably, the output of each register changes according to the input sequence and clock cycles, and the initial value of each register in a simulation step is taken from the output of the previous simulation. These transitions reflect the activity of circuit nodes, which is the primary source of dynamic power consumption. Therefore, transition probabilities provide a quantitative measure of the dynamic switching characteristics of the circuit, and can be used as an indicator for dynamic power analysis.

### v2

@@ 3.2. Function Shift Learning on Predictive Task
By ignoring the dependencies among input, i.e. by assuming they are independent, we can derive the local function, which is formally stated in Definition 3.2. Although this approximation is computationally efficient with an $O(1)$ complexity, it fails to capture the true function of the circuit.
@@
Building on these properties, we propose to learn the function shift: $y_{\phi}^{FSL}=y_{\phi}^{global}-y_{\phi}^{local},$ which measures the discrepancy between the local function (Definition 3.2) and the true global function (Definition 3.1). This formulation decouples the global function into two components: a simple local function and a function shift. Rather than predicting the complex global function directly, our model is trained to predict only the function shift. This isolates the complex contextual effects caused by reconvergence, allowing the global function to be reconstructed by simply combining the predicted shift with the local function.
@@
(5) $\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\Big[\big|\psi(x_{i})-y_{i}^{FSL}\big|\Big],$
where $x_{i}$ is the embedding for node $i$ and $\psi(\cdot)$ is a regression head.
@@
7:   $\hat{y}_{v}^{global}\leftarrow\hat{y}_{v}^{FSL}+\underbrace{\phi_{v}(\hat{y}_{u_{1}}^{global},\dots,\hat{y}_{u_{|\mathcal{N}(v)|}}^{global})}_{\text{Local Function}}$
8:Return $\hat{\mathbf{y}}^{global}$ for all nodes
Inference Stage.
3.3. Contrastive Task
Contrastive learning is a standard self-supervised strategy for learning representations of circuit functionality (Wang et al., 2024; Fang et al., 2025b, c; Wu et al., 2025). The fundamental principle is to pull embeddings of functionally equivalent circuits closer together in the embedding space while pushing apart those of functionally different circuits. This process trains the model to identify and encode the discriminative features that define a circuit’s intrinsic properties, all without needing explicit labels. Following prior work, we form training instances for each circuit $\mathcal{G}$. A positive sample, $\mathcal{G}^{+}$, is created by applying a functionally equivalent transformation to $\mathcal{G}$. All other circuits within the same batch are treated as negative samples, $\mathcal{G}^{-}$. We then optimize the encoder using the InfoNCE loss (Oord et al., 2018):
(6) $\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\;\mathcal{L}_{\text{InfoNCE}}(\mathcal{G},\mathcal{G}^{+},\mathcal{G}^{-}).$
The objective is to structure the embedding space so that functional equivalence corresponds to proximity, mapping similar circuits to nearby points while separating from non-equivalent ones.
Table 1. Comparison of contrastive task across various modalities(%).
@@ 3.3. Contrastive Task
4. Experiment
4.1. Implement Details
Dataset
In this paper, we conduct experiments on three modalities: RTL, AIG and PM netlist. For RTL, we follow prior works (Fang et al., 2025b, c) and collect data from ITC (Corno et al., 2002) and OpenCores (Albrecht, 2005). For combinational AIGs, we use ForgeEDA (Shi et al., 2025a). For Sequential AIGs, we follow DeepSeq2 (Khan et al., 2025) and extract sub-circuits from ITC (Corno et al., 2002), OpenCores (Albrecht, 2005) and ISCAS’89 (Brglez et al., 1989). For PM netlist, we follow prior work (Shi et al., 2025b) and extract sub-circuits from ForgeEDA (Shi et al., 2025a) and DeepCircuitX (Li et al., 2025). We summarize the statistics of the datasets used in both contrastive tasks (Table 2) and predictive tasks (Table 3). Since the global function in Definition 3.1 is hard to acquire, we follow previous works (Shi et al., 2023, 2025b) and approximate it using random simulations.
Table 2. Dataset for contrastive tasks with $\{min.,avg.,max.\}$.
RTL AIG PM Netlist
#Graphs 1138 67706 67728
#Nodes { 9, 102.6, 1987} {22, 162.2, 2127} {16, 88.6, 1192}
#Edges {10, 278.3, 2645 } {22, 176.0, 2275} {20, 110.5, 1361}
Depth {2, 8.9, 27} {6, 16.7, 29} {2, 5.9, 12}
Table 3. Dataset for predictive tasks with $\{min.,avg.,max.\}$.
Com. AIG Seq. AIG PM Netlist
#Graphs 9800 10007 83042
#Nodes {10, 2324.9, 45409} {23, 236.5, 1281} {12, 668.1, 4783}
#Edges {12, 3268.57, 68676} {21, 260.5, 1915} {11, 1113.9, 8914}
Depth {4, 49.1, 2657} {4, 21.0, 102} {1, 16.0, 260}
Evaluation Metrics
In this work, we evaluate our model on two types of tasks: contrastive and predictive. For the first type, the contrastive task, the goal is retrieval. Given a query circuit, the model must identify its functionally equivalent (positive) counterpart from a pool of $N$ candidate circuits. For our experiments, we set the pool size to $N=256$ for RTL and $N=1024$ for AIG and PM Netlists. We measure performance using the Recall@k (Rec@k) metric, reporting scores for $k\in\{1,5,10\}$. For the second type, the predictive tasks, we assess the model’s ability to determine node-level properties. We follow previous works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2025b) and perform logic-1 probability prediction, similarity prediction and transition probability prediction. We evaluate the accuracy of these predictions using Mean Absolute Error (MAE) and the $R^{2}$ score.
4.2. Contrastive Tasks
As shown in Table 1, TRACE consistently outperforms all baselines across the RTL, AIG, and PM Netlist modalities. Traditional message-passing GNNs (GCN, GIN, etc.) and specialized encoders (CircuitEncoder, FGNN2, DeepCell) achieve moderate performance, with Rec@1 scores generally ranging from 80% to 88.7%. However, Graph Transformer variants (GraphGPS, SGFormer, DIFFormer) show highly variable results: while competitive on RTL (up to 88.28%), they suffer severe performance degradation on AIGs and PM Netlists (Rec@1 dropping as low as 15.4% or below 50%) and can face OOM errors, highlighting their difficulty in handling circuit graph hierarchy. In contrast, TRACE achieves superior and stable results, pushing Rec@1 to 94.45% (RTL), 92.68% (AIGs), and 90.81% (PM Netlists), with Rec@10 consistently to 99.9%, 99.5% and 99.4% respectively. These results, which outperforms the second-best method by a large margin on RTL, AIG, and PM Netlist modalities, demonstrates its strong generalization capacity.
Table 4. Comparison of predictive tasks on combinational and sequential AIGs.
@@ 4.2. Contrastive Tasks
Table 5. Comparison of predictive task on PM netlists.
@@
Metric Message Passing Neural Network Graph Transformer TRACE
GCN GraphSAGE GAT GIN DeepCell GraphGPS SGformer DIFFormer
R2 0.718 0.946 0.902 0.734 0.942 0.846 0.918 0.696 0.994
MAE 0.112 0.048 0.059 0.102 0.053 0.083 0.056 0.141 0.013
4.3. Predictive Tasks
As shown in Table 4 and Table 5, TRACE demonstrates a clear and consistent superiority across all predictive tasks by achieving minimal MAE and best $R^{2}$, validating its strong generalization. On Combinational AIGs (Table 4), where traditional message-passing baselines and Graph Transformers struggle, TRACE delivers outstanding performance. TRACE achieves an MAE of 0.015 for logic-1 probability prediction and significantly outperforms all baselines in similarity prediction with an MAE of 0.055. The performance gap is further emphasized on Sequential AIGs, whose temporal dependencies pose additional challenges. Here, TRACE achieves substantial gains, reaching a minimal MAE of 0.009 for logic-1 probability prediction and an MAE of 0.005 for transition probability prediction, demonstrating superior precision. Finally, on PM Netlists (Table 5), TRACE again surpasses all baselines, achieving a minimal MAE of 0.013. This consistent ability to minimize error across modalities, in contrast to the varied performance of other models, emphasizes the adaptability of our approach. Overall, by consistently delivering the minimal MAE scores, TRACE demonstrates its robust capacity to generalize across diverse combinational and sequential circuit graph domains.
Figure 4. Error Accumulation Analysis. We illustrate the MAE of logic-1 probability prediction across different logic levels. Levels containing fewer than 10 nodes are excluded for better visualization. Top: Results on AIGs. Bottom: Results on PM netlists.
Error Accumulation
The iterative inference procedure in Algorithm 1 raises potential concerns about error accumulation. We investigate this by analyzing performance stability across different logic levels. As shown in Figure 4, we plot the MAE for the logic-1 probability prediction task, grouping nodes by their logic level. The results demonstrate that TRACE maintains a stable and outstanding performance across all levels. This indicates the robustness of TRACE to potential error accumulation and generalizes effectively to circuit graphs of varying depths.
Table 6. Ablation study on function shift learning (FSL).
Task GraphSAGE DeepCell DeepGate4 TRACE
w/o w. w/o w. w/o w. w/o w.
AIG Prob. 0.143 0.043 - - 0.027 0.020 0.024 0.015
Sim. 0.078 0.071 - - 0.078 0.076 0.066 0.055
PM Netlist Prob. 0.048 0.046 0.053 0.047 - - 0.036 0.013
4.4. Ablation Study
We conducted an ablation study to quantify the contribution of our proposed Function Shift Learning (FSL) component. The results in Table 6, compare our model (TRACE), second-based synchronous MPNNs (GraphSAGE) and asynchronous MPNNs (DeepGate4, DeepCell) with (w.) and without (w/o) the FSL component.
The findings demonstrate that FSL is crucial for achieving high performance. Our model, TRACE with FSL, consistently outperforms its ablated counterpart across all tasks. The impact is particularly significant on the PM Netlist probability prediction, where FSL reduces the MAE from 0.036 to 0.013, a relative error reduction of 63.9%. For the AIG tasks, FSL also yields significant gains, reducing the probability prediction MAE by 37.5% (from 0.024 to 0.015) and the similarity prediction MAE by 16.7% (from 0.066 to 0.055).
While FSL provides a consistent performance boost for baseline MPNNs on AIGs, they gain only a slight boost on the PM Netlists modality. This suggests that standard MPNNs struggle to leverage FSL on circuit graphs where nodes represent more complex functions, highlighting their limitation in capturing both the global function the function shift—a gap TRACE is designed to address.
5. Conclusion
@@
In this work, we introduced TRACE, a new paradigm for learning on circuit graphs that addresses the architectural limitations of conventional MPNNs and Transformers. By employing a novel Hierarchical Transformer and a function shift learning objective, TRACE directly models the position-aware, hierarchical nature of computation. Our extensive experiments on electronic circuits demonstrate that TRACE substantially outperforms all prior architectures, establishing a new state of the art. This work provides a proof of principle for a more architecturally sound approach to learning on circuit graphs, offering a powerful framework with potential applications across various domains.